# Completeness of coalition logics with seriality, independence of agents, or determinism

### v1

@@ 3.4 Upward derivability lemma
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$ is $\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{IA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}$ $\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
@@ 3.5 Completeness by induction
    Assume the modal depth of $\phi$ is 00.
    The $\phi$ is a formula of the classical propositional logic. As ${\mathsf{MCL}+\mathsf{X}}$ extends the classical propositional logic, $\vdash_{\mathsf{MCL}+\mathsf{X}}\phi$.
@@
    Assume the modal depth of $\phi$ is $n$, which is greater than 00.
@@
  5. (5)
    Assume the modal degree of $\chi_{i}$ is $n$. Let $\chi_{i}=\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
    By the downward validity lemma, the $\mathsf{X}$-validity-reduction-condition of $\chi_{i}$ is met:
@@
      the following two conditions hold:
      1. i.
      2. ii.
    By the inductive hypothesis, the $\mathsf{X}$-derivability-reduction-condition of $\chi_{i}$ is met:
    1. (a)
      $\vdash_{{\mathsf{MCL}+\mathsf{X}}}\gamma$, or
    2. (b)
      the following two conditions hold:
      1. i.
      2. ii.
@@ 4 Further work
  • $\langle[\mathrm{A}]\rangle\mathsf{X}\phi$: some joint strategy of $\mathrm{A}$ ensures $\phi$ at the next moment.
  • $\langle[\mathrm{A}]\rangle\mathsf{G}\phi$: some joint strategy of $\mathrm{A}$ ensures $\phi$ at all moments in the future.
  • $\langle[\mathrm{A}]\rangle(\phi\mathsf{U}\psi)$: some joint strategy of $\mathrm{A}$ ensures $\phi$ until $\psi$.
$\mathsf{ATL}$ is a very useful logic. What are the logics which are temporal extensions of the other seven coalition logics? We will study this in the future.
@@ Acknowledgments
Thanks go to the audience of seminars at University of Toulouse and University of Bergen, and the Third International Workshop on Logic of Multi-agent Systems at Zhejiang University, especially Thomas Ågotnes, Rustam Galimullin, and Emiliano Lorini.

### v2

@@ 3.6 Upward derivability lemma
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$ is $\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{IA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}$ $\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
@@ 3.7 Completeness by induction
Assume the modal depth of $\phi$ is 0.
Then $\phi$ is a formula of the classical propositional logic. As $\mathsf{MCL}+\mathsf{X}$} extends the classical propositional logic, $\vdash_{\mathsf{MCL}+\mathsf{X}}\phi$.
Assume the modal depth of $\phi$ is $n$, which is greater than 0.
Assume the modal degree of $\chi_{i}$ is $n$. Let $\chi_{i}=\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
By the downward validity lemma, the $\mathsf{X}$-validity-reduction condition of $\chi_{i}$ is met:
@@
    the following two conditions hold:
    1. (a)
    2. (b)
By the inductive hypothesis, the $\mathsf{X}$-derivability-reduction condition of $\chi_{i}$ is met:
@@
    $\vdash_{{\mathsf{MCL}+\mathsf{X}}}\gamma$, or
@@
    the following two conditions hold:
@@ 4 Further work
  • $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\mathsf{X}\phi$: some joint strategy of $\mathrm{A}$ ensures $\phi$ at the next moment.
  • $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\mathsf{G}\phi$: some joint strategy of $\mathrm{A}$ ensures $\phi$ at all moments in the future.
  • $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\mathsf{U}\psi)$: some joint strategy of $\mathrm{A}$ ensures $\phi$ until $\psi$.
$\mathsf{ATL}$ is a very useful logic. What are the logics that are temporal extensions of the other seven coalition logics? We will study this in the future.
@@ Acknowledgments
Thanks go to the audience of seminars at the University of Toulouse and the University of Bergen, and the Third International Workshop on Logic of Multi-agent Systems at Zhejiang University, especially Thomas Ågotnes, Rustam Galimullin, and Emiliano Lorini.